Aqua Daily Report May 13, 2002 1130 GMT - May 14, 2002 1130 GMT

Status Report From: Goddard Space Flight Center
Posted: Tuesday, May 14, 2002

The following activities were completed:

  1. Uplinked the daily Master Command Loads and Ephemeris.
  2. Successfully completed the first Ascent Delta-V maneuver (burn start at 1214z, 335 sec duration). Gained valuable information regarding thruster duty cycles and steady state performance that will be applied to future ascent maneuvers. Steps completed over this reporting period:
    1. Delta V burn
    2. Yaw maneuver to return to nominal attitude
    3. Cat bed heaters off
    4. Re-enable FPM fault management
    5. Return AMSU A1 & A2 to full scan
    6. Disable Isovalve SCS
  3. A note on science data processing status. AMSU (scanning) and AIRS (non-scanning) data has been successfully received and processed to Level-1A products at the JPL DAAC. CERES data (with the covers closed) is being processed at the LaRC DAAC. Also, the NASDA EOC reports successful receipt of GBAD data from EDOS in support of AMSR/e.

The following activities are scheduled for the next 24 s:

  1. Perform an engineering test of Spacecraft 911.
  2. Power ON the HSB instrument, perform a functional test, and place the instrument in normal mode (there was an error in the previous daily, this activity has always been scheduled for later today).
  3. Load and dump SCSs in preparation for the GNC Sensor Calibrations that begin tonight. A negative 25 degree Yaw maneuver will be performed first (maneuver start 0143z), followed by a positive 25 degree Yaw maneuver (maneuver start 0457z). Pitch and Yaw maneuver “pairs” are scheduled for subsequent days.
  4. Load the CERES diagnostic memory patch to both the Fore and Aft instruments and dump the data.
  5. Release both the CERES Fore and Aft instrument brakes.

Aqua Spacecraft Status:

  1. Spacecraft State is currently Standby.
  2. GNC Mode is Fine Point Mode, HiFi Ephemeris.

Aqua Instrument Status:

  1. AIRS is in Decontaminate state, Quiet bus is ON, Noisy bus is ON, Cryocooler is OFF, Scan mirror rotating is OFF and earth shield is CLOSED.
  2. AMSR-E is in Modified Sleep mode, rotating at 4 RPM.
  3. AMSU A1 is in Full Scan mode.
  4. AMSU A2 is in Full Scan mode.
  5. CERES Aft is in Safe mode and the covers are CLOSED.
  6. CERES Fore is in Safe mode and the covers are CLOSED.
  7. HSB is in Survival mode and the antenna is in the TARGET position.
  8. MODIS is in Standby mode, Nadir Door is Unlatched and CLOSED, Solar Diffuser Door is Unlatched and CLOSED, and the Space View Door is Unlatched and CLOSED.

Activities Deferred:

  1. None.


  1. Nothing further to report on the apparent misalignment of approximately 0.2 degrees between Star Tracker Assembly (STA) 1 and STA2. The actions stated in the May 10 daily continue.
  2. There have been a total of 21 occurrences (3 new events uncovered yesterday from back-orbit data analysis) of the Command Telemetry Controller (CTC) receiving incomplete transfer frames from Transponder Interface Electronics (TIE) – (formerly CLCW errors not associated with commands). This anomaly has still not affected the ability to command the spacecraft and it has been safe at all times. The actions stated in the May 10 daily continue.
  3. A patch is being developed to address the timing issue between the CTC and the USO. Initial indications from testing with the ETSF were positive. The activities stated in the May 13 daily continue.

// end //

More status reports and news releases or top stories.

Please follow SpaceRef on Twitter and Like us on Facebook.